Ralph Henwood was paid a salary of $83,800 during 20-- by Odesto Company. In addition, during the year Henwood started his own business as a public accountant and reported a net business income of $65,000 on his income tax return for 20--. Compute the following:

a. The amount of FICA taxes that was withheld from his earnings during 20-- by Odesto Company.

OASDI $
HI $

b. Henwood's self-employment taxes on the income derived from the public accounting business for 20--.

OASDI $
HI $
Answer #1
a). Solution :- OASDI = 83800 * 6.20 % $5,195.60
HI = 83800 * 1.45 % $1,215.10
Conclusion :-
OASDI (Salary income) $5,195.60
HI on (Salary income) $1,215.10
b.
OASDI $ 55000*12.4%= 6820
HI $ 55000*2.9%= 1595